# Sequential Rank and Preference Learning with the Bayesian Mallows Model

### v1

@@ 4.5 Mixtures of Mallows Models
To test the conditional particle filter and the proposed algorithm’s ability to estimate the number of mixture components we randomly generated ten datasets with $m=5$ items, each having two mixture components with dispersion parameters $\bm{\alpha}=(0.3,0.6)^{\prime}$ and equal probabilities $\bm{\tau}=(0.5,0.5)^{\prime}$. Modal rankings were $\bm{\rho}=(1,2,3,4,5)^{\prime}$ and $\bm{\rho}=(5,4,3,2,1)^{\prime}$.
@@
Figure 6 shows how the logarithm of the marginal likelihood (16) of the models estimated for each dataset varied with the number of clusters. We see a clear "elbow" at the correct number of two clusters. In terms of Bayes factors, however, models with more than two clusters were preferred for all the simulated datasets, as can be seen by the fact that the marginal likelihood keeps increasing beyond the two-cluster solution. Note that the marginal likelihood is not readily available when estimating mixtures of Mallows models using MCMC algorithms. For example, Vitelli et al., (2017) and Crispino et al., (2019) selected the number of clusters based on within-cluster distances, which requires saving the distance to the cluster centroid for each user at each MCMC step.
Figure 6: Logarithm of the marginal likelihood as a function of the number of clusters for ten simulated datasets.
Figure 7: Posterior histograms of $\bm{\alpha}$ for the two-cluster solution, for each simulated dataset.
@@
For the two-cluster model, the average of the posterior mean of $\bm{\alpha}$ across the ten simulated datasets was $0.299$ for the smallest component and $0.591$ for the largest component, with probabilities $0.508$ and $0.492$, both very close to the values in the data generating distribution. The average posterior probability of the true modal ranking $\bm{\rho}_{c}$ in each mixture component was $0.855$ in the cluster with $\alpha_{c}=0.3$ and $1.000$ in the cluster with $\alpha_{c}=0.6$, confirming that the modal ranking is easier to identify with a higher precision parameter.
@@ 5 Discussion
We have proposed an SMC2 algorithm for sequential estimation of the Bayesian Mallows model. The algorithm naturally incorporates data in the form of partial rankings and both transitive and non-transitive pairwise preferences, and is straightforward to parallelize. Compared to MCMC, the algorithm is competitive in use cases where data arrive sequentially and the posteriors of interest need to be recomputed for each new data batch. In batch estimation problems, MCMC is typically faster, but an advantage of SMC2 is that the marginal likelihood is readily available at no additional computational cost, which is useful for model comparison.
A number of future extensions are possible. First, conditioning on the full particle history in the conditional particle filter may lead to a high degree of degeneracy (Whiteley,, 2010), with the consequence that a large number of particle filters is required to obtain a sufficiently accurate approximation of the conditional posterior. Backward sampling has been shown to considerably decrease this degeneracy (Lindsten and Schön,, 2012), and would be interesting to consider in the current setting. Another extension of practical interest is to allow users to provide updated rankings, e.g., comparisons of previously unseen items. Stein, (2023, Ch. 6) proposed an algorithm for this in the case of partial rankings, in which users whose new data contradicted their current latent rankings were removed from the pool and then reentered. An extension of this approach to SMC2 would likely require a particle filter which performs this type of correction while still yielding unbiased estimated of the marginal likelihood $p(\bm{y}_{\mathcal{I}_{1:t}}|\theta)$, as equation (10) does in the current case where new users arrive at each timepoint. Finally, it may be of interest to let some or all of the parameters in $\theta$ to depend on time, e.g., to monitor how preferences in a population evolve. An MCMC algorithm for time-varying modal ranking $\bm{\rho}$ has been proposed by Asfaw et al., (2017), but sequential estimation is likely a good alternative in this case.

### v2

@@ 5.4 Mixtures of Mallows Models
To test the conditional particle filter and the proposed algorithm’s ability to estimate the number of mixture components we randomly generated ten datasets with $m=5$ items and $N=200$ users and ten datasets with $m=5$ items and $N=1000$ users, each having two mixture components with dispersion parameters $\bm{\alpha}=(0.3,0.6)^{\prime}$ and equal probabilities $\bm{\tau}=(0.5,0.5)^{\prime}$. Modal rankings were $\bm{\rho}=(1,2,3,4,5)^{\prime}$ and $\bm{\rho}=(5,4,3,2,1)^{\prime}$.
@@
Figure 7 shows how the logarithm of the marginal likelihood (16) of the models estimated for each dataset varied with the number of clusters. For both the $N=200$ case and the $N=1000$ case we see a clear "elbow" at the correct number of two clusters. In terms of Bayes factors, however, for the $N=200$ case models with more than two clusters were preferred for all the simulated datasets, as can be seen by the fact that the marginal likelihood keeps increasing beyond the two-cluster solution. For the $N=1000$ case, on the other hand, the marginal likelihood is almost completely flat beyond the two-cluster solution. This suggests that this estimator has too high variance in the $N=200$ case. Note that the marginal likelihood is not readily available when estimating mixtures of Mallows models using MCMC algorithms. For example, Vitelli et al., (2017) and Crispino et al., (2019) selected the number of clusters based on finding an "elbow" in a plot of within-cluster distances versus the number of clusters. This method, which requires saving the distance to the cluster centroid for each user at each MCMC step, can also be used as an alternative cluster selection method with SMC.
Figure 7: Logarithm of the marginal likelihood as a function of the number of clusters for ten simulated datasets with 200 users (left) and 1000 users (right).
Figure 8: Posterior histograms of $\bm{\alpha}$ for the two-cluster solution, for each simulated dataset with 200 users.
@@
For the two-cluster model with $N=200$, the average of the posterior mean of $\bm{\alpha}$ across the ten simulated datasets was $0.316$ for the smallest component and $0.587$ for the largest component, with probabilities $0.502$ and $0.498$, both very close to the values in the data generating distribution. The average posterior probability of the true modal ranking $\bm{\rho}_{c}$ in each mixture component was $0.855$ in the cluster with $\alpha_{c}=0.3$ and $1.000$ in the cluster with $\alpha_{c}=0.6$, confirming that the modal ranking is easier to identify with a higher precision parameter. With $N=1000$, the average posterior mean of $\bm{\alpha}$ was $(0.303,0.612)^{\prime}$, with probabilities $(0.500,0.500)^{\prime}$. The average posterior probability of the true modal ranking was $1.000$ for each component.
5.5 Timing Comparisons for Sequential Estimation
To compare the relative speed of SMC2 and MCMC for sequential estimation, we performed an additional set of experiments in which a set of users were assumed to arrive one at a time, and the posterior had to be updated after the arrival of each new user. For MCMC, this requires rerunning the whole algorithm at each timepoint, whereas for SMC2 a single iteration of Algorithm 3 is sufficient.
For a total number of timepoints equal to 10, 30, 50, 70, and 90 we generated 100 random datasets with $m=10$ items and $\alpha=0.1$. In addition to the complete rankings case, we also did experiments in which only the top-5 or the top-3 items were retained. For each simulated dataset, 5,000 samples from the posterior distribution were obtained. For MCMC we used a burn-in of 1000 and no thinning, so that a total of 6,000 iterations were run each time a new observation arrives. For SMC2, the number of particles was set to 5,000 which were either processed on a single core or in parallel on eight cores. The initial number of particle filters was set to 1, and the thresholds for resampling and particle filter doubling were set identically to Section 5.2. We emphasize that it is hard to ensure that the effective posterior sample size obtained from SMC2 and MCMC are identical and hence obtain decisive evidence of which algorithm is faster. Our goal here was rather to understand the scaling behavior of the respective algorithms for sequential estimation, and how this depends on the amount of missing data.
Figure 9: Wall time for sequential estimation with complete rankings (left), top-5 rankings (center), and top-3 rankings (right). Shaded areas are 95% confidence intervals.
The results are shown in Figure 9. In all cases, MCMC scaled approximately quadratically. On the other hand, SMC on a single core and on eight cores scaled close to linearly with complete rankings. In the case with missing rankings, SMC scaled quadratically, but remained faster than MCMC in the top-5 case. In the top-3 case, however, the computing time for SMC2 grew more quickly with the number of users than MCMC. Closer examination of the data revealed that this was due to the particle filter doubling required to properly integrate over the latent variables in these cases.
6 Sequential Analysis of Formula 1 Data
We now analyze the Formula 1 data introduced in Figure 1. As mentioned in the introduction, the rankings were derived from the results of each race. Whenever a driver either was disqualified, did not finish a race, or did not start, the ranking for the particular race was set to missing and assumed to have a higher value than all the ranked drivers. This yielded a missingness proportion of 12.7%. Each race $n\in\{1,\dots,68\}$ was hence modeled as an assessor yielding a top-$k_{n}$ ranking, with $k_{n}$ equal to the number of drivers completing the race.
We computed the posterior distribution of the Bayesian Mallows model with footrule distance sequentially using the SMC2 algorithm with $R=10^{5}$ particles and $S=10$ initial particle filters. We used multinomial resampling, uniform latent rank proposal, resampling threshold $A=R/2$, and threshold $B=0.2$ for particle filter doubling. We confirmed that the number of particle filters was sufficiently large by running the model multiple times with different random number seeds and checking that the posterior quantities of interest remained essentially the same. At the final timepoint, the posterior mean of the scale parameter $\alpha$ was 0.170 with 95% posterior interval $(0.165,0.178)$.
Figure 10: CP consensus over time for Formula 1 drivers. The black lines show the CP consensus ranking, and the color scale shows the posterior probability of having the given ranking or higher.
At each timepoint we computed the cumulative probability (CP) consensus based on the current posterior distribution. The CP consensus (Vitelli et al.,, 2017, Sec. 5.1) was computed by first selecting the driver with the highest posterior probability of being ranked first. Then among the remaining drivers we found the driver with the highest posterior probability of being ranked first or second, and so on. The resulting trace plots are shown in Figure 10, in which the solid black lines indicate the CP consensus ranking of each driver and the color scale indicates the posterior probability. Note that by construction, the posterior probability for the driver whose CP consensus equals 16 – the number of items – always has probability 1, since all drivers must be ranked 16 or higher. Also note that the true ranking $\bm{\rho}$ is assumed to be constant, and that Figure 10 shows how our posterior belief about this parameter develops as we obtain more data.
Some interesting features can be seen from Figure 10. First we note that the CP consensus changes quickly during the first ten races, as more data is incorporated into the posterior and the effect of the uniform prior fades. Correspondingly, the probability for the drivers ranked as best after the first rounds is relatively low. For example Charles Leclerc is ranked first in the first eight timepoints, with a probability between 0.16 and 0.76 and Carlos Sainz is ranked second during the first three timepoints, with a probability between 0.29 and 0.65. After the first ten races the consensus stabilizes. For example, Max Verstappen is ranked first for the first time after the eighth race with a probability of 0.91 and retains this rank until the last timepoint with a probability exceeding 0.99 from the thirteenth timepoint and onward. However, there is also movement in the later timepoints; for example, Charles Leclerc climbs from rank 8 at timepoint 54 to rank 4 at the final timepoint.
@@ 7 Discussion
We have proposed an SMC2 algorithm for sequential estimation of the Bayesian Mallows model. The algorithm naturally incorporates data in the form of partial rankings and both transitive and non-transitive pairwise preferences, and is straightforward to parallelize. Compared to MCMC, the algorithm is competitive in use cases where data arrive sequentially and the posteriors of interest need to be recomputed for each new data batch. In batch estimation problems, on the other hand, MCMC has been faster in all cases considered in this paper.
A number of future extensions are possible. First, conditioning on the full particle history in the conditional particle filter may lead to a high degree of degeneracy (Whiteley,, 2010), with the consequence that a large number of particle filters is required to obtain a sufficiently accurate approximation of the conditional posterior. Backward sampling has been shown to considerably decrease this degeneracy (Lindsten and Schön,, 2012), and would be interesting to consider in the current setting. Next, a challenge with SMC2 is that its computational complexity grows quadratically with time, as the plot for top-3 rankings in Figure 9 (right) illustrates. A related nested particle filter algorithm which scales linearly with time has been proposed by Crisan and Míguez, (2018). Since their algorithm requires the parameters to be real-valued, it is not directly applicable to the Mallows model, but creating such an extension is an interesting problem for future research. Another extension of practical interest is to allow users to provide updated rankings, e.g., comparisons of previously unseen items. Stein, (2023, Ch. 6) proposed an algorithm for this in the case of partial rankings, in which users whose new data contradicted their current latent rankings were removed from the pool and then reentered. An extension of this approach to SMC2 would likely require a particle filter which performs this type of correction while still yielding unbiased estimates of the marginal likelihood $p(\bm{y}_{\mathcal{I}_{1:t}}|\theta)$, as equation (10) does in the current case where new users arrive at each timepoint. Finally, it may be of interest to let some or all of the parameters in $\theta$ depend on time, e.g., to monitor how preferences in a population evolve. An MCMC algorithm for time-varying modal ranking $\bm{\rho}$ has been proposed by Asfaw et al., (2017), but sequential estimation is likely a good alternative in this case. More recently Piancastelli and Barreto-Souza, (2025) have proposed a model for timeseries of rankings, based on the Mallows model. In this model, complete or partial rankings of a set of items are assumed to be observed on a relatively large number of timepoints, and rather than estimating the consensus ranking the focus is on timeseries parameters describing the dynamics by which the modal ranking changes over time.